suppose A=3i-j-2k and B=2i+3j+k,Find |AxB|,(A+2B)x(2A-B),(A+B)x(A-B)
1
Expert's answer
2020-04-24T19:07:42-0400

Lets write the cross product formula:

AxB ="\begin{vmatrix}
 i & j & k\\
 A_1 & A_2 & A_3\\
 B_1 & B_2 & B_3
\end{vmatrix} ="

= (A2B3 - A3B2)i - (A1B3 - A3B1)j + (A1B2 - A2B1)k


The length of vector is:

"|a|=\sqrt{a_1^2+a_2^2+a_3^2}"


The sum of two vectors is:

a + b = {a1 + b1; a2 + b2; a3 + b3}


The difference of two vectors is:

a - b = {a1 - b1; a2 - b2; a3 - b3}


So let's find |AxB|:

AxB = "\begin{vmatrix}
 i & j & k\\
 3 & -1 & -2\\
 2 & 3 & 1
\end{vmatrix} ="

= ((-1)*1 - (-2)*3)i - (3*1 - (-2)*2) + (3·3 - (-1)·2)k = (-1 + 6)i - (3 + 4)j + (9 + 2)k = {5; -7; 11}

|AxB| = "\sqrt{5^2+(-7)^2+11^2}=\sqrt{25+49+121}=\sqrt{195}\approx" 13.96


Now let's find (A+2B)x(2A-B):

2B = {2*B1; 2*B2; 2*B3} = {2*2; 2*3; 2*1} = {4; 6; 2}

2A = {2*A1; 2*A2; 2*A3} = {2*3; 2*(-1); 2*(-2)} = {6; -2; -4}

A + 2B = {3 + 4; (-1) + 6; (-2)+2} = {7; 5; 0}

2A - B = {6 - 2; (-2) - 3; (-4) - 1} = {4; -5; -5}

(A + 2B)x(2A - B) = "\begin{vmatrix}
 i & j & k\\
 7 & 5 & 0\\
 4 & -5 & -5
\end{vmatrix} ="

= (5*(-5) - 0*(-5))i - (7*(-5) - 0*4)j + (7*(-5) - 5*4)k = (-25 - 0)i - (-35 - 0)j + (-35 - 20)k = {-25; 35; -55}


Finally let's find (A + B)x(A - B):

A + B = {3 + 2; (-1) + 3; (-2) + 1} = {5; 2; -1}

A - B = {3 - 2; (-1) - 3; (-2) - 1} = {1; -4; -3}

(A + B)x(A - B) = "\begin{vmatrix}
 i & j & k\\
 5 & 2 & -1\\
 1 & -4 & -3
\end{vmatrix} ="

= (2*(-3) - (-1)*(-4))i - (5*(-3) - (-1)*1)j + (5*(-4) - 2*1)k = (-6 - 4)i - (-15 + 1)j + (-20 - 2)k = {-10; 14; -22}